how to resolve force F=100N and angle of 30 degrees into its horizontal and vertical component
1
Expert's answer
2019-08-21T05:01:01-0400

if the angle to the horizontal:


"F_H=F*\cos(30°)\approx88.6 N"

Horizontal component


"F_V=F*\sin(30°)=50N"

Vertical component;

if the angle to the vertical:


"F_H=F*\sin(30°)=50 N
\\F_V=F*\cos(30°)\approx88.6N"